### 29 Feb 2024 v1

@@ Abstract
We present the All-Seeing Project V2: a new model and dataset designed for understanding object relations in images. Specifically, we propose the All-Seeing Model V2 (ASMv2) that integrates the formulation of text generation, object localization, and relation comprehension into a relation conversation (ReC) task. Leveraging this unified task, our model excels not only in perceiving and recognizing all objects within the image but also in grasping the intricate relation graph between them, diminishing the relation hallucination often encountered by Multi-modal Large Language Models (MLLMs). To facilitate training and evaluation of MLLMs in relation understanding, we created the first high-quality ReC dataset (AS-V2) which is aligned with the format of standard instruction tuning data. In addition, we design a new benchmark, termed Circular-based Relation Probing Evaluation (CRPE) for comprehensively evaluating the relation comprehension capabilities of MLLMs. Notably, our ASMv2 achieves an overall accuracy of 52.04 on this relation-aware benchmark, surpassing the 43.14 of LLaVA-1.5 by a large margin. We hope that our work can inspire more future research and contribute to the evolution towards artificial general intelligence. Our project is released at https://github.com/OpenGVLab/all-seeing.
@@ 1 Introduction
(3) To evaluate the relation comprehension ability of existing MLLMs, we construct the CRPE benchmark. Notably, our ASMv2 achieves an overall accuracy of 52.04 on CRPE, surpassing the 43.14 of LLaVA-1.5 by a large margin. We also evaluate ASMv2 on various image-level and region-level vision-language tasks. Specifically, our model achieves an overall score of 74.4 on MMBench [51] and 1621.0 on MME [18], surpassing LLaVA-1.5 [49] by 5.5 points and 90.0 points separately. Besides, the average accuracy of ASMv2 on grounding benchmarks [30, 57] is 87.42, outperforming Qwen-VL [2] by 1.69 points.
@@ 5.3 Results on Open-ended Scene Graph Generation
In this section, we evaluate the Relation Conversation capability of our model through the Open-ended Scene Graph Generation task on the Panoptic Scene Graph (PSG) dataset [84], which is a widely-used benchmark for the scene graph generation. See Sec. 0.C.1 for the results on the Predicate Classification task.
@@ 5.4 Results on CRPE
Qwen-VL [2] 76.34 19.47 26.99 35.36 27.27
LLaVA-1.5 [49] 84.90 37.86 43.68 47.89 43.14
ASMv2 (ours) 87.24 50.55 48.46 57.10 52.04
@@
As shown in Tab. 6, the performance of existing MLLMs on the Existence questions is significantly higher than on the Subject, Predicate, and Object questions. This suggests that these models have a more robust capability to recognize objects within an image than to comprehend the relations between them. Specifically, our ASMv2 shows a remarkable improvement in understanding object relations compared to the other models. For example, ASMv2 achieves an overall accuracy of 52.04, which is significantly higher than the 43.14 of LLaVA-1.5 and the 27.27 of Qwen-VL. These results demonstrate that our can comprehend the relations between the objects within the image better, benefiting from the training of relation conversation data.
@@ Appendix 0.A Relation Conversation
Figure 6: A complex example for the formulation of Relation Conversation.
@@
An example is shown in Fig. 6. As discussed in Sec. 4.1, to parse this example into a scene graph, we first assign the semantic tag “people” to the bounding box highlighted in red and blue. Similarly, we assign the semantic tag “grass” to the bounding box highlighted in green. We then extract the predicate label enclosed in “<pred></pred>” (i.e., “standing on”) and the box coordinates of its subjects and objects (i.e., bounding boxes highlighted with bold underline). After that, we utilize these box coordinates as keys to match their respective semantic tags. Considering that two subjects are linked to the predicate while only one object is linked, we broadcast the object to match the number of subjects. Given $N$ subjects and $N$ objects, we pack them into $N$ tuples in order, where each tuple consists of one subject and one object. In this example, we obtain two tuples, resulting in two parsed triplet (people, standing on, grass), each connected to a people with different bounding boxes.
@@ Appendix 0.B The All-Seeing Dataset v2
More data examples of AS-V2 are shown in Figs. 7, 8 and 9. Besides, prompts used to generate detailed description data are shown in Tabs. 7 and 8
@@ Appendix 0.C The All-Seeing Model v2
0.C.1 Predicate Classification
@@
0.C.2 Implementation Details
Task #Samples Dataset
Captioning 124K TextCaps [71], ShareGPT4V [9]
VQAv2 [20], GQA [24], OKVQA [58], A-OKVQA [67],
VQA 314K ScienceQA [56], CLEVR [29], Visual7W [106]
OCR 157K ST-VQA [4], LLaVAR [99], OCR-VQA [59], DocVQA [12]
Grounding 643K RefCOCO/+/g [30, 57]
RegionVQA 2.3M RefCOCOg [57], VG [31], VCR [91], AS-Core [80]
Text 40K ShareGPT [68]
0.C.2.1 Training Stage 1.
0.C.2.2 Training Stage 2.
The global batch size is set to 512 and the learning rate is initialized as $2\times 10^{-3}$ in both the pre-training phase and the instruction-tuning phase. The language model and vision-language connector are trainable in both phases while the vision encoder is always frozen. We train the model for 5000 steps in the pre-training phase and 1 epoch in the instruction-tuning phase. The other settings remain the same as the instruction-tuning phase of Stage 1.

### 21 Mar 2024 v2

@@ Abstract
We present the All-Seeing Project V2: a new model and dataset designed for understanding object relations in images. Specifically, we propose the All-Seeing Model V2 (ASMv2) that integrates the formulation of text generation, object localization, and relation comprehension into a relation conversation (ReC) task. Leveraging this unified task, our model excels not only in perceiving and recognizing all objects within the image but also in grasping the intricate relation graph between them, diminishing the relation hallucination often encountered by Multi-modal Large Language Models (MLLMs). To facilitate training and evaluation of MLLMs in relation understanding, we created the first high-quality ReC dataset (AS-V2) which is aligned with the format of standard instruction tuning data. In addition, we design a new benchmark, termed Circular-based Relation Probing Evaluation (CRPE) for comprehensively evaluating the relation comprehension capabilities of MLLMs. Notably, our ASMv2 achieves an overall accuracy of 64.50 on this relation-aware benchmark, surpassing the 55.63 of LLaVA-1.5 by a large margin. We hope that our work can inspire more future research and contribute to the evolution towards artificial general intelligence. Our project is released at https://github.com/OpenGVLab/all-seeing.
@@ 1 Introduction
(3) To evaluate the relation comprehension ability of existing MLLMs, we construct the CRPE benchmark. Notably, our ASMv2 achieves an overall accuracy of 64.50 on CRPE, surpassing the 55.63 of LLaVA-1.5 by a large margin. We also evaluate ASMv2 on various image-level and region-level vision-language tasks. Specifically, our model achieves an overall score of 74.4 on MMBench [51] and 1621.0 on MME [18], surpassing LLaVA-1.5 [49] by 5.5 points and 90.0 points separately. Besides, the average accuracy of ASMv2 on grounding benchmarks [30, 57] is 87.42, outperforming Qwen-VL [2] by 1.69 points.
@@ 5.3 Results on Open-ended Scene Graph Generation
In this section, we evaluate the Relation Conversation capability of our model through the Open-ended Scene Graph Generation task on the Panoptic Scene Graph (PSG) dataset [84], which is a widely-used benchmark for the scene graph generation. See Sec. 0.C.2 for the results on the Predicate Classification task.
@@ 5.4 Results on CRPE
Qwen-VL [2] 85.11 45.66 38.19 31.60 38.48
LLaVA-1.5 [49] 88.69 57.44 54.24 55.21 55.63
ASMv2 (ours) 92.14 69.21 58.95 65.34 64.50
@@
As shown in Tab. 6, the performance of existing MLLMs on the Existence questions is significantly higher than on the Subject, Predicate, and Object questions. This suggests that these models have a more robust capability to recognize objects within an image than to comprehend the relations between them. Specifically, our ASMv2 shows a remarkable improvement in understanding object relations compared to the other models. For example, ASMv2 achieves an overall accuracy of 64.50, which is significantly higher than the 55.63 of LLaVA-1.5 and the 38.48 of Qwen-VL. These results demonstrate that our model can comprehend the relations between the objects within the image better, benefiting from the training of relation conversation data.
@@ Appendix 0.A Relation Conversation
As shown in Fig. 6, to parse this example into a scene graph, we first assign the semantic tag “people” to the bounding box highlighted in red and blue. Similarly, we assign the semantic tag “grass” to the bounding box highlighted in green. We then extract the predicate label enclosed in “<pred></pred>” (i.e., “standing on”) and the box coordinates of its subjects and objects (i.e., bounding boxes highlighted with bold underline). After that, we utilize these box coordinates as keys to match their respective semantic tags. Considering that two subjects are linked to the predicate while only one object is linked, we broadcast the object to match the number of subjects. Given $N$ subjects and $N$ objects, we pack them into $N$ tuples in order, where each tuple consists of one subject and one object. In this example, we obtain two tuples, resulting in two parsed triplet (people, standing on, grass), each connected to a people with different bounding boxes.
@@ Appendix 0.B The All-Seeing Dataset v2
More data examples of AS-V2 are shown in Figs. 7, 8 and 9. Besides, prompts used to generate detailed description data are shown in Tabs. 7 and 8.
@@ Appendix 0.C The All-Seeing Model v2
0.C.1 Implementation Details
Task #Samples Dataset
Captioning 124K TextCaps [71], ShareGPT4V [9]
VQAv2 [20], GQA [24], OKVQA [58], A-OKVQA [67],
VQA 314K ScienceQA [56], CLEVR [29], Visual7W [106]
OCR 157K ST-VQA [4], LLaVAR [99], OCR-VQA [59], DocVQA [12]
Grounding 643K RefCOCO/+/g [30, 57]
RegionVQA 2.3M RefCOCOg [57], VG [31], VCR [91], AS-Core [80]
Text 40K ShareGPT [68]
0.C.1.1 Training Stage 1.
0.C.1.2 Training Stage 2.
The global batch size is set to 512 and the learning rate is initialized as $2\times 10^{-5}$ in both the pre-training phase and the instruction-tuning phase. The language model and vision-language connector are trainable in both phases while the vision encoder is always frozen. We train the model for 5000 steps in the pre-training phase and 1 epoch in the instruction-tuning phase. The other settings remain the same as the instruction-tuning phase of Stage 1.
0.C.2 Predicate Classification
@@
0.C.3 Ablation Study
Table 11: Ablation results. We report the $Q\rightarrow AR$ accuracy for VCR [91] and the overall accuracy for CRPE. REC denotes the average accuracy score across RefCOCO [30], RefCOCO+ [57], and RefCOCOg [57].
Ablation Settings MME MM-Vet REC VCR CRPE
ASMv2 1621.0 41.3 87.4 78.4 64.5
- Stage1 Training 1527.6 35.6 86.1 78.7 64.7
- Relation Conversation 1554.6 42.2 86.6 77.0 55.3
In this section, we ablate the training settings of ASMv2. The experimental settings are the same as those discussed in Sec. 0.C.1. As shown in Tab. 11, the two-stage training process of ASMv2 and the utilization of relation conversation data is essential for achieving excellent performance on both image-level and region-level benchmarks simultaneously. We can observe that skipping the first training stage leads to significant performance degradation on MME [18], MM-Vet [90], and Referring Expression Comprehension (REC) Benchmarks [30, 57], indicating that the model struggles to understand visual information at both the image and region levels simultaneously without the two-stage training strategy.
Furthermore, removing relation conversation data from the training corpora results in inferior performance on REC [30, 57], VCR [91], and CRPE. The performance on MME [18] also experiences a drop of about 66.4 points, primarily due to a decrease in the count score, from 170.0 to 155.0, and a decrease in the position score, from 163.3 to 133.3. These results demonstrate the effectiveness of our relation conversation data in improving capabilities for region-level visual information understanding and relation comprehension.